# Science Opportunities of Wet Extreme Mass-Ratio Inspirals

### v1

@@ II.5 Cosmology with Wet EMRIs
Wet EMRIs provide a promising avenue for probing cosmological models by serving as both “bright” and “dark” sirens. These systems combine GW measurements of the luminosity distance with EM observations to determine the redshift of their possible host galaxies. When the host galaxy can be unambiguously identified, wet EMRIs act as bright sirens, enabling an independent measurement of the Hubble parameter. However, to qualify as a bright siren, a wet EMRI requires high localization precision and a unique association with its host AGN.
To assess the feasibility of identifying bright sirens, we analyze catalogs of wet EMRIs generated from various astrophysical models (see Table 1), such as an $\alpha$-disk model with $T_{\rm disk}=10^{6}$ years. For each EMRI system, Fisher analysis is employed to estimate the number of potential host galaxies within its localization volume. The solid angle uncertainty, $\Delta\Omega_{s}$, derived from the Fisher matrix components associated with the angular parameters $(\theta_{s},\phi_{s})$, is expressed as (see also [75, 122]):
where $\Sigma_{ij}=(\Gamma^{-1})_{ij}$ is the covariance matrix obtained as the inverse of the Fisher information matrix $\Gamma_{ij}$. This covariance matrix quantifies the uncertainties and correlations in the angular parameters. To estimate the volume of three-dimensional localization ($V_{\text{sky}}$) in spherical coordinate, the uncertainty of the solid angle is combined with the uncertainty of the luminosity distance $d_{L}$. The localization volume is given by
$V_{\text{sky}}=\frac{1}{2}r^{3}(z)\,\sigma(\ln{d_{L}})\,\Delta\Omega_{s}\,,$ (25)
where $r(z)$ is the comoving distance to the source as a function of redshift, and $\sigma(\ln{d_{L}})$ is the relative uncertainty in the distance measurement. For LISA-detected EMRIs, Fisher analysis indicates that $V_{\text{sky}}$ can be restricted to $\mathcal{O}(10^{2})\,\rm Mpc^{3}$ at moderate redshifts ($z<0.3$), depending on the SNR and the inclusion of higher order waveform harmonics, demonstrating the high localization precision of LISA.
After applying a conservative AGN fraction ($f_{\rm AGN}=1\%$), we operationally classify an EMRI as a bright siren if the expected number of AGNs within this volume is less than $1.1$. Across various models in Table 1, we find that approximately $20\%\sim 25\%$ of detectable wet EMRIs have resolvable host galaxies, translating to a detection rate of $3\sim 30$ bright sirens per year. However, not all wet EMRIs meet this criterion. Those systems with multiple potential AGN hosts within the localization volume are classified as dark sirens, indicating ambiguity in host galaxy association.
@@ II.5.1 Wet EMRIs as bright sirens
To measure $H_{0}$ using wet EMRI bright sirens, the luminosity distance $d_{\rm L}$ obtained from GW signals is combined with the redshift $z$ of the host AGN. Assuming a flat $\Lambda$CDM cosmology, the relationship between luminosity distance and redshift ($d_{\rm L}-z$ relation) is expressed as
@@ II.5.2 Wet EMRIs as dark sirens
[…]
@@
This method offers a promising pathway for advancing GW cosmology, particularly in scenarios where EM counterparts are absent. Furthermore, the combination of bright and dark sirens provides complementary constraints across different redshift ranges: bright sirens probe lower redshifts with uniquely identifiable hosts, while dark sirens extend measurements to higher redshifts using statistical associations. Together, these approaches not only enhance our ability to constrain $H_{0}$ and other cosmological parameters but also offer valuable insights into the “Hubble tension”—a significant discrepancy between early-universe $H_{0}$ measurements from the Cosmic Microwave Background (CMB)[128, 129] and late-universe measurements from standard candles like Cepheid-calibrated Type Ia supernovae[124, 130] or GW-EM observations [123, 131, 132]. This tension, which challenges the standard $\Lambda$CDM model, has spurred exploration of potential solutions, including modified gravity [133, 134, 135, 136, 137, 138, 139, 140], alternative forms of dark energy [141, 142, 143, 144, 145], or unaddressed measurement systematics [146, 147, 148, 149, 150]. By combining the complementary strengths of bright and dark sirens, GW cosmology provides a robust and independent approach to addressing this critical cosmological puzzle, while also offering a more comprehensive understanding of the universe’s expansion history.

### v2

@@ II.5 Cosmology with Wet EMRIs
Wet EMRIs provide a promising avenue for probing cosmological models by serving as both “bright” and “dark” sirens. These systems combine GW measurements of the luminosity distance with EM observations to determine the redshift of their possible host galaxies. When the host galaxy can be unambiguously identified, wet EMRIs act as bright sirens, enabling an independent measurement of the Hubble parameter. However, to qualify as a bright siren, a wet EMRI requires high localization precision and a unique association with its host AGN. In practice, selection effects including galaxy survey incompleteness and AGN luminosity function need to be taken into account in a self-consistent way, e.g., a Bayesian approach that has been widely used for population inference in previous GW cosmology works [142, 143]. As a proof of concept, we idealize the host identification by assuming a complete galaxy survey where galaxies are either detectable AGNs or others (non-detectable AGNs or normal galaxies). With this simplification, the host AGN identification is established as long as the EMRI error volume is sufficiently small containing only one AGN. As shown by forecast in Sec. II.1.2, only low-redshift ($z\lesssim 0.3$) host AGNs can be identified with EMRI GW signals. In the following forecast of host AGN identification, we adopt a conservative assumption that $1\%$ of low-redshift galaxies are detectable AGNs (see Sec. II.3 for a brief summary of AGN detection).
@@ II.5.1 Wet EMRIs as bright sirens
[…]
@@
The inclusion of wet EMRI bright sirens in cosmological studies enhances the scope of GW-based standard siren measurements, complementing those from massive black hole binaries (MBHBs), binary neutron stars, and binary black holes. In parallel, MBHBs detected by space-based interferometers like LISA also serve as powerful bright sirens. As demonstrated by Tamanini et al. [146], LISA can constrain $H_{0}$ to sub-percent precision, particularly in its most optimistic configurations, with uncertainties reaching as low $0.5\%$. Importantly, MBHB standard sirens extend to redshifts as high as $z\sim 8$, significantly beyond the range accessible to wet EMRIs and EM probes. This broader redshift coverage allows MBHBs to provide independent and complementary constraints on cosmological parameters.
@@ II.5.2 Wet EMRIs as dark sirens
$p(D_{i}\,|\,\mathbf{\Theta},\mathcal{H},I)=$ $\int\int\,dz_{\rm gw}\,dd_{L}\,p(d_{L}\,|\,z_{\rm gw},\mathbf{\Theta},\mathcal{H},I)\,$
@@
This method offers a promising pathway for advancing GW cosmology, particularly in scenarios where EM counterparts are absent. Furthermore, the combination of bright and dark sirens provides complementary constraints across different redshift ranges: bright sirens probe lower redshifts with uniquely identifiable hosts, while dark sirens extend measurements to higher redshifts using statistical associations. By combining the complementary strengths of bright and dark sirens, GW cosmology offers a robust and independent means of investigating the universe’s expansion history, leading to a more comprehensive understanding of its evolution.